OpenStudy (anonymous):

How many liters of fluorine gas can react with 32.0 grams of sodium metal at standard temperature and pressure? Show all of the work used to find your answer. 2Na + F2 yields 2NaF Can anyone plz help me out with this? @hartnn @Abhisar @Callisto @paki @Preetha

OpenStudy (abhisar):

brb

OpenStudy (anonymous):

okay @Abhisar

OpenStudy (abhisar):

m back :D

OpenStudy (abhisar):

ok..now temme the atomic mass of sodium

OpenStudy (anonymous):

22.98977

OpenStudy (abhisar):

yes..take it 23

OpenStudy (anonymous):

okay :)

OpenStudy (abhisar):

Now from the equation we can see that 2 moles of sodium needs 1 mole of fluorine gas. So 23*2=46 grams of sodium needs one mole of fluorine

OpenStudy (abhisar):

So 32 grams will need 32/46 = 0.69 moles of fluorine

OpenStudy (abhisar):

Now, one mole of any gas at STP has a volume of 22.4 litrres, so 0.69 moles of fluorine will have 0.69*22.4=15.58 L

OpenStudy (anonymous):

and that will be our answer?

OpenStudy (abhisar):

15.58 L

OpenStudy (abhisar):

ok..lso y don't u try to put it urself on the basis of ur understanding and i'll correct u

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

@Abhisar

OpenStudy (abhisar):

no of moles of fluorine that reacts with 32 gram of Na = 32/46=0.6956 1 mole of Fluorine at STP=22.4 L So 0.69 moles of fluorine = 22.4*0.6956=15.58 L

OpenStudy (anonymous):

Okay got it!! thanks sooo much for all your help your awesome!

OpenStudy (abhisar):

\(\rlap{\color{blue}{\huge\bigstar}\huge \color{blue}{ \text{You're Most Welcome! }}\color{blue}\bigstar}{\; \color{aqua}{\huge\bigstar}\huge \color{aqua}{\text{You're Most Welcome! }}\color{aqua}\bigstar}\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)